Testbank Problem 148

The following information was taken from the books and records of Ivanhoe, Inc.:

1. Net Income $409,200

2. Capital structure:

a. Convertible 6% bonds. Each of the 280, $1,000 bonds is convertible into 50 shares of common stock at the present date and for the next 10 years. 280,000

b. $10 par common stock, 220,000 shares issued and outstanding during the entire year. 2,200,000

c. Stock warrants outstanding to buy 14,560 shares of common stock at $20 per share.

3. Other information: a. Bonds converted during the year None

b. Income tax rate 30%

c. Convertible debt was outstanding the entire year

d. Average market price per share of common stock during the year $32

e. Warrants were outstanding the entire year

f. Warrants exercised during the year None

Compute basic earnings per share. (Round answer to 2 decimal places, e.g. 52.75.)

Basic earnings per share $

Compute diluted earnings per share. (Round answers to 2 decimal places, e.g. 52.75.)

Security EPS

Common Stock $

Warrants $

Conversion Bonds $
Answer #1

Net Income = $409,200

Number of shares = 220,000 shares

Basic EPS = $ 1.86

Security Net Income Adjustment Adjusted Net Income Shares Adjustment Adjustment shares Diluted EPS
Common Stock 409,200 - 409,200 220,000 - 220,000 1.86
Warrants 409,200 - 409,200 220,000 5460 [note 2] 231060 1.77
Convertible Bonds 409,200 11.760 [note 1] 420960 225600 14000 239600 1.76

Notes:

1. $2,80,000*0.06*0.70 = $11.760

2. Warrants = [14560*20 ]/ 32 =  291200/32 = 9100

  14,560 - 9100 = 5460